Calculate the percent available chlorine in chlori
Answer #1

“Available chlorine” is calculated as if Cl2 was present as the oxidizing agent. This can be done, because both Cl2 and ClO- react with the iodide ion, I-, in a 1:1 stoichiometry as shown in the equations below. Cl2(g) + 2 I-(aq)® 2 Cl-(aq) + I2(aq) ClO-(aq) + 2 I-(aq) + H2O(l) ® I2(aq) + Cl-(aq) + 2 OH-(aq) Available chlorine is reported as the mass of “chlorine” per unit mass of solution, or powder if this is a dry chlorine bleach. This mass ratio is frequently converted to a percent and reported as the percent “available chlorine.”Ans-

Available chlorine is calculated as if  Cl2 was present as the oxidizing agent.  This can be done, because both Cl2 and ClO- react with the iodide ion, I-, in a 1:1 stoichiometry as shown in the equations below.

Cl2(g) + 2 I-(aq)®  2 Cl-(aq) + I2(aq)

ClO-(aq) + 2 I-(aq) + H2O(l) ® I2(aq) + Cl-(aq) +  2 OH-(aq)

Available chlorine is reported as the mass of chlorine per unit mass of solution, or powder if this is a dry chlorine bleach. This mass ratio is frequently converted to a percent and reported as the percent available chlorine.

Chlorine is soluble in water and its dissolution in water is a physical and chemical process. 2, 4 Upon contact with water, chlorine disproportionates , giving chloride ion and hypochlorous acid 2, 8 or hypochlorite ion, depending on the pH of the solution. Thus, dissolved chlorine is present in aqueou 2 s solutions as a mixture of three active species, molecular chlorine (Cl ), hypochlorous acid (HOCl), and hypochlorite ion (OCl ). These three species are generally called free available chlorine – 9 (FAC). The distribution of these species can be described by the following equations 8 Cl2 2 H + H O ¾ Cl + H + HOCl K = 4×10 (25°C)

–4 HOC a l ¾ H + OCl pK = 7.54 (25°C) (2) + –

Both equilibria are pH dependent. Thus, the distribution of the three species varies with the pH of the solutions. By using the equilibrium constants,